OpenStudy (anonymous):

Tips on finding the appropriate equations for a transformation "T" using given bounds for a region R in the xy-plane?

OpenStudy (anonymous):

I'm at a loss on how to do this...an example from the book is: R is bounded by y=2x-1, y=2x+1, y=1-x, y=3-x. Find equations for a transformation T that maps a rectangular region S in the uv-plane onto R, where the sides of S are parallel to the u- and v-axes.

OpenStudy (perl):

is this related to calculus

OpenStudy (anonymous):

yes - part of "change of variables in multiple integrals"

OpenStudy (anonymous):

calc 3

OpenStudy (perl):

first lets graph the region R https://www.desmos.com/calculator/nhjjlqitio

OpenStudy (perl):

thats a parallelogram, and we want to turn it into a rectangle

OpenStudy (anonymous):

yes, understood

OpenStudy (perl):

y=2x-1, y=2x+1, y=1-x, y=3-x Let's rewrite these questions y-2x = -1 y -2x =1 y+x = 1 y + x= 3 this suggests the change of variables u = y -2x v = y + x

OpenStudy (anonymous):

wow, that was much simpler than I thought it would be

OpenStudy (anonymous):

finding the change of variables, that is

OpenStudy (perl):

Now the lines in R correspond to u = -1, u = 1 , v = 1, v = 3

OpenStudy (perl):

in the uv plane

OpenStudy (perl):

we still have to find the transformation T

OpenStudy (anonymous):

Yes, the change of variables was really all I needed help on. It seemed so arbitrary and didn't really know what to represent u and v with

OpenStudy (perl):

oh ok :)

OpenStudy (anonymous):

thank you so much! :)